THIS IS A CONVERSION FACTORS QUESTION

Camp nirvana sent a warning message over the CB radio. The glass skulls are patrolling the roadways and only allowing free passage with a payment of 100 kg glass. The latest report stated there will be a possible run in every 400 miles. Your team has some solid 30 lb glass blocks to use as bribe “money”. For a trip of 2695 miles, how many blocks should you take on your trip to ensure you have enough glass?

if you guys could answer my question that will be great please and thank you. :)
1
Expert's answer
2020-11-12T06:36:51-0500

no. of bribes needed = "\frac{2695}{400} = 6.74 = 7 bribes"


1 bribe = 100kg = 100(2.21 Ib) = 221Ib

"\therefore" Amount per bribe = 221Ib per bribe


Amount of glass used for bribe = 221Ib per bribe × 7bribes = 1547 Ib


1glass block = 30 Ib

x glass blocks = 1547Ib


x = "\frac{1547}{30}" = 51.6 glass blocks



"\therefore" you should take about 52 glass blocks with you on your trip to ensure that you have enough glass.